# Robust and Efficient Zeroth-Order LLM Fine-Tuning via Adaptive Bayesian Subspace Optimizer

### v2

@@ 1 Introduction
    We propose BSZO, a zeroth-order optimizer that uses Bayesian inference to aggregate gradient information across multiple perturbation directions within a subspace. To our knowledge, this is the first application of Kalman filtering to ZO optimization for LLMs.
@@ 2 Related Work
Bayesian methods provide a principled way to integrate observations with prior knowledge while quantifying uncertainty. Kalman filtering (Kalman, 1960) is the canonical example: it sequentially updates a Gaussian belief over a hidden state as new measurements arrive. Gaussian processes extend this idea to function-space modeling and underpin Bayesian optimization (Shahriari et al., 2016; Williams and Rasmussen, 2006). Our work adapts the Kalman perspective to ZO gradient estimation: we model the projected gradient as a hidden state, interpret each perturbation query as a noisy linear measurement, and update a posterior that pools information across all sampled directions within an iteration. Leveraging the flexibility of the Bayesian framework, we further design an adaptive residual mechanism that effectively fuses both historical and current-batch information. This yields improved gradient estimates without additional memory overhead.
@@ 3.3 Posterior Update In Subspace
According to the standard Bayesian linear regression theory (Williams and Rasmussen, 2006), after $m$ perturbations and observations $(d^{(1)},\hat{y}^{(1)}),\ldots,(d^{(m)},\hat{y}^{(m)})$, the posterior $\tilde{g}|Y\sim\mathcal{N}(\mu^{(m)},\Sigma^{(m)})$ is also a Gaussian distribution, where
@@ 3.4 Algorithm
Algorithm 1 Bayesian Subspace Zeroth-Order Optimization (BSZO)
for $t=1$ to $T$ do
  Sample $k$ random seeds $\{s_{i}\}_{i=1}^{k}$
  Initialize cache $Y\leftarrow\{\}$
  for $\tau=1$ to $m$ do
   if $\tau\leq k$ then
    $d\leftarrow e_{\tau}$
    $y\leftarrow(\mathcal{L}(\theta)-f_{0})/\varepsilon$
    $Y[\tau]\leftarrow y$ $\triangleright$ Cache directional derivative
   else
    $d\leftarrow e_{j}$
    $y\leftarrow Y[j]$ $\triangleright$ Reuse cached value
   end if
   $K\leftarrow\Sigma d/(d^{\top}\Sigma d+\sigma_{e}^{2})$
  end for
  for $i=1$ to $k$ do
  end for
end for
return $\theta$
$\textsc{Randn}(n,s)$: returns $n$-dim Gaussian vector seeded by $s$
@@ 4 Theoretical Analysis
Definition 4.1.
$\sigma_{e}^{2}=\sigma_{\varepsilon}^{2}+\text{tr}(\Sigma),$ (14)
Theorem 4.2.
Corollary 4.3.
where $\Delta_{0}:=\mathcal{L}(\theta_{0})$-$\mathcal{L}^{*}$.
5 Experiments
5.1 Experimental Setup
5.2 Performance in Masked Language Models
BSZO achieves stable and competitive performance on mLMs. As shown in Table 2, BSZO-B reaches 78.01% average accuracy on RoBERTa-large, surpassing MeZO (76.10%, +1.91%), MeZO-Adam (75.15%, +2.86%), HiZOO (69.05%, +8.96%), and LOZO (74.73%, +3.28%). BSZO achieves 77.30% average accuracy, second only to BSZO-B. BSZO secures top results on SST-2 (92.89%) and WIC (57.68%), while BSZO-B excels on TREC (85.8%). Both variants demonstrate strong and consistent performance across all tasks.
Table 2: Test accuracy (%) on RoBERTa-large (355M). We show the test accuracy of MeZO, MeZO-Adam, HiZOO, LOZO, BSZO, and BSZO-B, with the top two results highlighted in bold. BSZO-B is the baseline version of BSZO.
Method SST-2 RTE CB WIC TREC Avg
MeZO 92.43 69.68 87.50 49.69 81.2 76.10
MeZO-Adam 92.32 62.09 82.14 54.39 84.8 75.15
HiZOO 91.97 59.21 78.57 50.31 65.2 69.05
LOZO 91.63 59.57 87.50 54.55 80.4 74.73
BSZO 92.89 68.59 83.93 57.68 83.4 77.30
BSZO-B 92.66 67.87 85.71 57.99 85.8 78.01
@@ 5.3 Performance in decoder-only models
5.4 Memory Usage
@@ 5.5 Ablation Study
Table 4 shows ablation results on OPT-1.3B for two design choices of BSZO: subspace dimension $k$ and sample count $m$. In Table 4(a), when $m=k$, RTE accuracy climbs from 60.29% ($k=1$) to 67.51% ($k=4$), while SST-2 peaks at $k=8$ (93.23%), suggesting that increasing $k$ generally improves performance. In Table 4(b), with extra refinement ($m=k+1$), RTE performance improves consistently. Comparing to Table 4(a), $m=k+1$ boosts RTE by 1-2% at most $k$ levels (e.g., from 64.26% to 66.79% at $k=2$, from 66.07% to 68.59% at $k=8$). This confirms that the adaptive sampling step refines the posterior estimate. For SST-2, performance improves notably at $k=4$ but shows little gain at other $k$ values. This suggests that adaptive sampling has varied effects across tasks.
Table 4: Ablation studies on OPT-1.3B. (a) Effect of subspace dimension $k$ with $m=k$. (b) Effect of observation count $m$ with $m=k+1$. (c) Noise-free adaptive sampling. Best per row in bold.
(a) Effect of $k$ (b) Effect of $m$ (c) NF-Adaptive
$k$ SST-2 RTE $k$ SST-2 RTE $k$ SST-2 RTE
1 92.32 60.29 1 91.74 61.37 1 91.28 63.58
2 92.78 64.26 2 92.43 66.79 2 92.43 65.34
4 92.66 67.51 4 93.58 66.43 4 93.12 66.07
8 93.23 66.07 8 93.23 68.59 8 93.35 69.31
@@ 6 Conclusion
In this work, we introduce BSZO, which is the first zeroth-order optimizer that applies Kalman filtering to aggregate gradient information across multiple perturbation directions for LLM fine-tuning. By treating finite-difference measurements as noisy observations of the true gradient, BSZO builds a posterior distribution over the projected gradient and refines it through Bayesian updates. We design a residual-based adaptive mechanism to adjust the perturbation scale adaptively without manual tuning. Our theoretical analysis shows that BSZO improves the convergence rate by a factor of $k/\gamma$ over standard ZO methods. Experiments on RoBERTa, Mistral, and OPT show that BSZO obtains competitive accuracy across various tasks while keeping memory usage close to inference-only baselines.

### v3

@@ 1 Introduction
    We propose BSZO, a zeroth-order optimizer that uses Bayesian inference to aggregate gradient information across multiple perturbation directions within a subspace. To our knowledge, this is the first application of Bayesian inference and Kalman filtering to ZO optimization for LLMs.
@@ 2 Related Work
Bayesian methods provide a principled way to integrate observations with prior knowledge while quantifying uncertainty. Kalman filtering (Kalman, 1960) is the canonical example: it sequentially updates a Gaussian belief over a hidden state as new measurements arrive. Gaussian processes extend this idea to function-space modeling and underpin Bayesian optimization (Shahriari et al., 2016; Rasmussen and Williams, 2006). Our work adapts the Kalman perspective to ZO gradient estimation: we model the projected gradient as a hidden state, interpret each perturbation query as a noisy linear measurement, and update a posterior that pools information across all sampled directions within an iteration. Leveraging the flexibility of the Bayesian framework, we further design an adaptive residual mechanism that effectively fuses both historical and current-batch information. This yields improved gradient estimates without additional memory overhead.
@@ 3 Method
[…]
Method SST-2 RTE CB WIC TREC Avg
MeZO 92.22 ($\pm$0.42) 66.35 ($\pm$3.06) 86.07 ($\pm$5.56) 55.20 ($\pm$3.73) 85.36 ($\pm$2.33) 77.04
MeZO-Adam 92.34 ($\pm$0.50) 63.61 ($\pm$1.41) 81.07 ($\pm$2.71) 52.85 ($\pm$4.19) 78.80 ($\pm$5.76) 73.73
HiZOO 91.44 ($\pm$0.45) 59.21 ($\pm$2.46) 76.43 ($\pm$1.96) 53.60 ($\pm$2.93) 63.44 ($\pm$2.61) 68.82
LOZO 91.83 ($\pm$0.30) 62.60 ($\pm$2.31) 84.29 ($\pm$3.87) 54.20 ($\pm$1.32) 77.76 ($\pm$2.15) 74.14
BSZO 92.66 ($\pm$0.21) 67.80 ($\pm$1.52) 85.71 ($\pm$1.79) 56.05 ($\pm$1.47) 84.16 ($\pm$0.54) 77.28
BSZO-B 92.27 ($\pm$0.41) 68.38 ($\pm$1.94) 84.29 ($\pm$1.49) 57.21 ($\pm$0.98) 84.80 ($\pm$1.57) 77.39
Algorithm 1 Bayesian Subspace Zeroth-Order Optimization (BSZO)
for $t=1$ to $T$ do
  Sample $k$ random seeds $\{s_{i}\}_{i=1}^{k}$
  Initialize cache $Y\leftarrow\{\}$
  for $\tau=1$ to $m$ do
   if $\tau\leq k$ then
    $d\leftarrow e_{\tau}$
    $y\leftarrow(\mathcal{L}(\theta)-f_{0})/\varepsilon$
    $Y[\tau]\leftarrow y$ $\triangleright$ Cache directional derivative
   else
    $d\leftarrow e_{j}$
    $y\leftarrow Y[j]$ $\triangleright$ Reuse cached value
   end if
   $K\leftarrow\Sigma d/(d^{\top}\Sigma d+\sigma_{e}^{2})$
  end for
  for $i=1$ to $k$ do
  end for
end for
return $\theta$
$\textsc{Randn}(n,s)$: returns $n$-dim Gaussian vector seeded by $s$
@@ 3.3 Posterior Update In Subspace
According to the standard Bayesian linear regression theory (Rasmussen and Williams, 2006), after $m$ perturbations and observations $(d^{(1)},\hat{y}^{(1)}),\ldots,(d^{(m)},\hat{y}^{(m)})$, the posterior $\tilde{g}|Y\sim\mathcal{N}(\mu^{(m)},\Sigma^{(m)})$ is also a Gaussian distribution, where
@@ 4 Theoretical Analysis
Definition 4.1.
$\sigma_{e}^{2}=\sigma_{\varepsilon}^{2}+\text{tr}(\Sigma),$ (14)
Theorem 4.2.
Corollary 4.3.
where $\Delta_{0}:=\mathcal{L}(\theta_{0})$-$\mathcal{L}^{*}$.
5 Experiments
5.1 Experimental Setup
5.2 Performance in Masked Language Models
BSZO achieves stable and competitive performance on mLMs. As shown in Table 1, BSZO-B reaches 77.39% average accuracy on RoBERTa-large, surpassing MeZO (77.04%, +0.35%), MeZO-Adam (73.73%, +3.66%), HiZOO (68.82%, +8.57%), and LOZO (74.14%, +3.25%). BSZO achieves 77.28% average accuracy, second only to BSZO-B. BSZO secures top result on SST-2 (92.66%), while BSZO-B excels on RTE (68.38%) and WIC (57.21%). Moreover, BSZO exhibits notably lower variance across tasks (see Table 11 for raw results). Both variants demonstrate strong and consistent performance across all tasks.
@@ 5.3 Performance in decoder-only models
5.4 Memory and Time Efficiency
Table 4: Memory usage (GB) and per-step time (ms) across different models.
OPT-1.3B Mistral-7B OPT-13B
Method Mem Time Mem Time Mem Time
MeZO 9.1 109.7 18.3 283.9 30.0 464.2
MeZO-Adam 19.7 135.1 47.6 373.1 82.1 614.5
HiZOO 15.7 188.0 34.3 540.2 58.9 877.1
LOZO 9.3 102.0 18.3 274.2 30.0 452.0
BSZO 9.8 97.0 18.8 275.7 30.1 440.5
@@ 5.5 Ablation Study
Table 5: Ablation studies. (a) Effect of subspace dimension $k$ with $m=k$ on OPT-1.3B. (b) Effect of $m=k+1$ on OPT-1.3B. (c) Effect of adaptive noise on OPT-1.3B (bf16). (d) Effect of adaptive noise on OPT-13B (bf16). Best results in bold. Full results in fp32 are in Table 12.
(a) Effect of $k$ (b) Effect of $m$ (c) Adaptive Noise
$k$ SST-2 RTE $k$ SST-2 RTE $k$ w/ w/o
1 92.32 60.29 1 91.74 61.37 1 54.15 55.24
2 92.78 64.26 2 92.43 66.79 2 57.76 56.32
4 92.66 67.51 4 93.58 66.43 4 61.73 56.32
8 93.23 66.07 8 93.23 68.59 8 66.43 57.76
(d) Adaptive Noise on OPT-13B (bf16)
Method SST-2 RTE WSC COPA TREC WIC
w/ adaptive 93.23 69.31 61.54 83.00 79.20 56.27
w/o adaptive 91.97 63.18 58.65 85.00 75.00 54.70
Table 5 shows ablation results on OPT-1.3B for two design choices of BSZO: subspace dimension $k$ and sample count $m$. In Table 5(a), when $m=k$, RTE accuracy climbs from 60.29% ($k=1$) to 67.51% ($k=4$), while SST-2 peaks at $k=8$ (93.23%), suggesting that increasing $k$ generally improves performance. In Table 5(b), with extra refinement ($m=k+1$), RTE performance improves consistently. Comparing to Table 5(a), $m=k+1$ boosts RTE by 1-2% at most $k$ levels (e.g., from 64.26% to 66.79% at $k=2$, from 66.07% to 68.59% at $k=8$). This confirms that the adaptive sampling step refines the posterior estimate (see Table 12 for more details). Table 5(c) investigates adaptive noise under bf16 precision on OPT-1.3B. As $k$ grows, the gap between w/ and w/o adaptive noise becomes more pronounced: at $k=8$, the adaptive variant leads by 8.67% on RTE, indicating that adaptive noise yields substantial gains in low-precision settings. Table 5(d) validates this on OPT-13B (bf16), where adaptive noise brings improvements on 5 out of 6 tasks, with RTE gaining 6.13%.
@@ 6 Conclusion
In this work, we introduce BSZO, which is the first zeroth-order optimizer that applies Kalman filtering to aggregate gradient information across multiple perturbation directions for LLM fine-tuning. By treating finite-difference measurements as noisy observations of the true gradient, BSZO builds a posterior distribution over the projected gradient and refines it through Bayesian updates. We design a residual-based adaptive mechanism to adjust the perturbation scale adaptively without manual tuning. Our theoretical analysis shows that BSZO improves the convergence rate by a factor of $k/\gamma$ over standard ZO methods. Experiments on RoBERTa, Mistral, and OPT show that BSZO achieves strong accuracy across various tasks, remains stable under fp16/bf16 precision where existing methods often collapse, and keeps memory usage close to inference-only baselines.
@@ Appendix C Experiment Details
Appendix D Raw Experimental Results
We provide the complete raw results of 5 independent runs for each method on RoBERTa-large in Table 11. The mean and standard deviation reported in Table 1 are computed from these results.
Table 11: Raw test accuracy (%) of 5 runs on RoBERTa-large (355M).
Dataset Method Run 1 Run 2 Run 3 Run 4 Run 5
SST-2 MeZO 92.43 92.32 91.74 92.78 91.86
MeZO-Adam 92.32 92.66 91.51 92.43 92.78
HiZOO 91.97 91.86 91.28 91.17 90.94
LOZO 91.63 92.09 91.74 91.51 92.20
BSZO 92.89 92.43 92.78 92.43 92.78
BSZO-B 92.66 91.74 92.32 91.97 92.66
RTE MeZO 69.68 68.95 65.70 65.34 62.09
MeZO-Adam 62.09 64.26 64.62 64.98 62.09
HiZOO 59.21 63.18 57.76 56.68 59.21
LOZO 59.57 63.90 61.01 65.34 63.18
BSZO 68.59 68.23 69.68 66.07 66.43
BSZO-B 67.87 70.04 70.76 66.43 66.79
CB MeZO 87.50 85.71 91.07 76.79 89.29
MeZO-Adam 82.14 83.93 80.36 82.14 76.79
HiZOO 78.57 75.00 75.00 78.57 75.00
LOZO 87.50 82.14 82.14 89.29 80.36
BSZO 83.93 85.71 87.50 87.50 83.93
BSZO-B 85.71 83.93 82.14 85.71 83.93
WIC MeZO 49.69 58.31 52.98 57.52 57.52
MeZO-Adam 54.39 51.10 54.70 46.55 57.52
HiZOO 50.31 54.39 51.10 54.70 57.52
LOZO 54.55 54.55 51.88 55.17 54.86
BSZO 57.68 57.52 55.64 54.70 54.70
BSZO-B 57.99 57.99 55.80 56.58 57.68
TREC MeZO 81.20 86.40 86.40 86.20 86.60
MeZO-Adam 84.80 74.20 71.40 83.00 80.60
HiZOO 65.20 65.20 65.20 62.20 59.40
LOZO 80.40 74.80 77.20 79.20 77.20
BSZO 83.40 84.60 84.40 83.80 84.60
BSZO-B 85.80 84.60 85.20 82.20 86.20
Table 12: Full ablation studies on OPT-1.3B (fp32). (a) Effect of subspace dimension $k$ with $m=k$. (b) Effect of observation count $m$ with $m=k+1$. (c) Noise-free adaptive sampling. Best per row in bold.
(a) Effect of $k$ (b) Effect of $m$ (c) NF-Adaptive
$k$ SST-2 RTE $k$ SST-2 RTE $k$ SST-2 RTE
1 92.32 60.29 1 91.74 61.37 1 91.28 63.58
2 92.78 64.26 2 92.43 66.79 2 92.43 65.34
4 92.66 67.51 4 93.58 66.43 4 93.12 66.07
8 93.23 66.07 8 93.23 68.59 8 93.35 69.31